# VITA-Audio: Fast Interleaved Cross-Modal Token Generation for Efficient Large Speech-Language Model

### v1

@@ 4.3 Evaluation on Fundamental Speech Competence
TTS We evaluate the TTS performance of VITA-Audio on Seed-TTS Seed-TTS and LibriTTS LibriTTS benchmarks. We use Whisper-Large-V3 Whisper and ParaformerParaformer to transcribe into text the generated English and Chinese speech, respectively. For all VITA-Audio models, we use the following prompt during inference.
Prompt for TTS task.
    "messages": [
        "role": "user",
[…]
We present the results of VITA-Audio at each stage in Table 5. The experiments demonstrate that VITA-Audio outperforms other open-source models with a similar number of parameters. Additionally, it should be noted that, despite using ten MCTP modules for accelerated inference, VITA-Audio’s TTS capabilities are largely preserved throughout the training process, further validating the effectiveness of the MCTP module in aligning text and audio.
ASR We evaluate the ASR performance of the four stages of VITA-Audio on WenetSpeech WenetSpeech , AIshell AISHELL_1 , LibriSpeech Librispeech and Fleurs Fleurs , and the results are reported in Table 6 and Table 7. We follow the original data splits for each benchmark. For all VITA-Audio models, we use the following prompt during inference.
Prompt for ASR task.
    "messages": [
        "role": "user",
[…]
        <|begin_of_audio|> audio_sequence <|end_of_audio|>"
The results for other works are partially reproduced from their respective original works for comparison. In these results, the output of VITA-Audio’s first stage (S1) consists of text tokens directly generated by the LLM, while the outputs of the second (S2), third (S3), and fourth (S4) stages are alternately generated by both the LLM and the MCTP module.
Although the ASR performance of VITA-Audio does not demonstrate significant advantages, the ASR capability in S3 remains largely consistent with that of S1, indicating that VITA-Audio does not establish a unidirectional text-to-audio mapping. Instead, it constructs an alignment relationship between the text and speech modalities, facilitating efficient mapping between them.
@@ 4.4 Evaluation of Latency
As mentioned in Section 3.4, VITA-Audio-Vanilla only uses the main model for output; VITA-Audio-Turbo uses both the main model and all the MCTP modules during each forward pass; and VITA-Audio-Speed and VITA-Audio-Balance progressively increase the number of MCTP modules used to ensure higher accuracy.
As shown in Table 8, we present a comparison of the time consumption for different inference modes and model sizes. Noted that the time consumption does not include the cost of the speech encoder and speech decoder. We observe that in VITA-Audio-Turbo, a speedup of approximately 5×5$\times$5 × is achieved across models ranging from $0.5$B to $72$B, greatly enhancing the output token throughput. VITA-Audio-Speed also achieves around 3×3$\times$3 × acceleration across various sizes, resulting in a desirable performance for real-time speech dialogue systems. For example, $72$ B VITA-Audio generates approximately $50$ tokens per second, which, excluding generated text tokens, corresponds to roughly three seconds of audio and associated text when using a $12.5$Hz speech tokenizer. This performance is sufficiently fast for human-computer interaction.
@@
We visualize the timeline of model decoding phrase in Fig. 5. The green marks denote the tokens generated by the main model, and the blue marks are the tokens generated by MCTP modules. We set the number of prefiil tokens to $32$. And Fig. 5 shows that VITA-Audio-Turbo completes the generation of the first audio chunk in about $50$ ms, while VITA-Audio-Vanilla requires about $220$ ms. VITA-Audio-Speed and VITA-Audio-Balance generate fewer audio tokens in the first forward and more text audio tokens in the following forward. Thus, they are slower than VITA-Audio-Turbo but still significantly faster than VITA-Audio-Vanilla.
Table 9: Generation time (ms) of the first audio segment under different inference modes in streaming inference. To enable more real-time speech generation, we progressively increase the number of steps in the flow matching model during streaming inference. The table shows the decoding time when the sampling step of the flow matching model is set to $1$.
Boost 39 53 151 243
Vanilla 39 236 151 426

### v2

@@ 4.3 Evaluation on Fundamental Speech Competence
TTS We evaluate the TTS performance of VITA-Audio on Seed-TTS Seed-TTS and LibriTTS LibriTTS benchmarks. We use Whisper-Large-V3 Whisper and ParaformerParaformer to transcribe into text the generated English and Chinese speech, respectively.
We present the results of VITA-Audio at each stage in Table 3. In these results, the output of VITA-Audio’s first stage (S1) consists of text tokens directly generated by the LLM, while the outputs of the second (S2), third (S3), and fourth (S4) stages are alternately generated by both the LLM and the MCTP module. The experiments demonstrate that VITA-Audio outperforms other open-source models with a similar number of parameters. Additionally, it should be noted that, despite using ten MCTP modules for accelerated inference, VITA-Audio’s TTS capabilities are largely preserved throughout the training process, further validating the effectiveness of the MCTP module in aligning text and audio.
ASR We evaluate the ASR performance of the four stages of VITA-Audio on WenetSpeech WenetSpeech , AIshell AISHELL_1 , LibriSpeech Librispeech , and Fleurs Fleurs , and a subset of the results are reported in Table 4. More detailed results can be found in Table E3 and Table E4 in the Appendix. The results for other works are partially reproduced from their respective original works for comparison.
It can be observed that VITA-Audio-plus-vanilla demonstrates highly competitive performance across various benchmarks. Moreover, VITA-Audio-plus-Boost achieves remarkably fast inference speed while still maintaining strong overall performance.
@@ 4.4 Evaluation of Latency
As mentioned in Section 3.4, VITA-Audio-Vanilla only uses the main model for output; VITA-Audio-Turbo uses both the main model and all the MCTP modules during each forward pass; and VITA-Audio-Boost and VITA-Audio-Balance progressively increase the number of MCTP modules used to ensure higher accuracy.
As shown in Table 5, we present a comparison of the time consumption for different inference modes and model sizes. Noted that the time consumption does not include the cost of the audio encoder and audio decoder. We observe that in VITA-Audio-Turbo, a speedup of approximately $5\times$ is achieved across models ranging from 0.50.5B to 7272B, greatly enhancing the output token throughput. VITA-Audio-Boost also achieves around $3\times$ acceleration across various sizes, resulting in a desirable performance for real-time speech dialogue systems. For example, $72$ B VITA-Audio generates approximately $40$ tokens per second, which, excluding generated text tokens, corresponds to roughly three seconds of audio and associated text when using a 12.512.5Hz speech tokenizer. This performance is sufficiently fast for human-computer interaction.
@@
We visualize the timeline of model decoding phrase in Fig. 5. The green marks denote the tokens generated by the main model, and the blue marks are the tokens generated by MCTP modules. We set the number of prefiil tokens to $32$. And Fig. 5 shows that VITA-Audio-Turbo completes the generation of the first audio chunk in about $50$ ms, while VITA-Audio-Vanilla requires about $220$ ms. VITA-Audio-Boost and VITA-Audio-Balance generate fewer audio tokens in the first forward and more text audio tokens in the following forward. Thus, they are slower than VITA-Audio-Turbo but still significantly faster than VITA-Audio-Vanilla.
@@ 5 Conclusion
Acknowledgments
This work is partially funded by National Natural Science Foundation of China (Grant No. 62506158 and No. 62441234), and CCF-Tencent Rhino-Bird Open Research Fund.
[…]
Supplementary Material
Appendix A Training Data
Appendix B Training Pipline
Stage 4: Supervised Fine-tuning. After the previous three training stages, VITA-Audio has acquired the ability to efficiently and accurately map text to audio. To enable speech-to-speech dialogue capability, we then conduct supervised fine-tuning using speech QA datasets while maintaining a small amount of TTS, ASR, and text-only data to ensure training stability. To optimize training effectiveness, different learning rates are used for the MCTP module and the main LLM. For the speech-to-speech data, we employ an interleaved output format. This design enforces the model to initiate audio token generation during the first forward pass, enabling synchronized decoding of the audio tokens rather than waiting until all text tokens have been generated.
Appendix C Limitations
While our approach enables efficient generation of audio tokens, the overall end-to-end latency remains above the theoretical lower bound, primarily due to the constrained generation speed of the audio decoder. Further improving the response speed of the audio decoder in end-to-end speech models is a worthwhile direction for future exploration.
Appendix D Data Format
Speech QA Interleaved Data Format
    "messages": [
        "role": "user",
        "content": "<|begin_of_audio|> audio_sequence_1 <|end_of_audio|>"
        "role": "assistant",
        "content": "text_sequence_1 <|begin_of_audio|>
        audio_sequence_2 <|end_of_audio|> text_sequence_2
        <|begin_of_audio|> audio_sequence_3 <|end_of_audio|>"
        "role": "user",
        "content": "<|begin_of_audio|> audio_sequence_4 <|end_of_audio|>"
        "role": "assistant",
        "content": "text_sequence_3 <|begin_of_audio|>
        audio_sequence_5 <|end_of_audio|> text_sequence_4
        <|begin_of_audio|> audio_sequence_6 <|end_of_audio|>"
Prompt for TTS task.
    "messages": [
        "role": "user",
[…]
Prompt for ASR task.
    "messages": [
        "role": "user",
[…]
        <|begin_of_audio|> audio_sequence <|end_of_audio|>"
Appendix E Figures and Tables
Table E1: Summary of datasets used in VITA-Audio for different stages.
Table E2: Comparison of model structures between VITA-Audio and VITA-Audio-Plus.
Table E3: Results on Automatic Speech Recognition (ASR) Benchmarks. “Sxx” denotes the $x$-th training stage. Compared to other methods, VITA-Audio is trained with open-source data only.
Datasets Model WER (%) $\downarrow$ LibriSpeech Librispeech test-cleantest-other Qwen2-Audio-base  Qwen2-Audio 1.744.04 Baichuan-Audio-base  Baichuan-Audio 3.02 — 6.04 Freeze-Omni  Freeze-Omni 3.82 — 9.79 VITA-1.5 (S3)  VITA-1.5 3.40 — 7.50 LUCY (S1)  LUCY 3.36 — 8.05 Step-Audio-chat  Step-Audio 3.19 — 10.67 Qwen2.5-Omni  Qwen2.5-Omni 2.37 — 4.21 VITA-Audio-Turbo 6.29 — 12.86 VITA-Audio-Vanilla 2.98 — 8.07 VITA-Audio-Plus-Boost 3.13 — 7.07 VITA-Audio-Plus-Vanilla (S1) 1.91 — 4.29 VITA-Audio-Plus-Vanilla 2.00 — 4.60 Fleurs Fleurs zhen Qwen2-Audio-base  Qwen2-Audio 3.63 — 5.20 Baichuan-Audio-base  Baichuan-Audio 4.15 — 8.07 Step-Audio-chat  Step-Audio 4.26 — 8.56 Qwen2.5-Omni  Qwen2.5-Omni 2.924.17 VITA-Audio-Plus-Vanilla 3.69 — 4.54
Table E4: Results on Automatic Speech Recognition (ASR) Benchmarks. “Sxx” denotes the $x$-th training stage. Compared to other methods, VITA-Audio is trained with open-source data only.
Datasets Model WER (%) $\downarrow$ AISHELL-1 AISHELL_1 Qwen2-Audio-base  Qwen2-Audio 1.52 Baichuan-Audio-base  Baichuan-Audio 1.93 Freeze-Omni  Freeze-Omni 2.48 LUCY (S1)  LUCY 2.40 Step-Audio-chat  Step-Audio 2.14 Qwen2.5-Omni  Qwen2.5-Omni 1.13 VITA-Audio-Turbo 7.70 VITA-Audio-Vanilla 4.46 VITA-Audio-Plus-Boost 4.72 VITA-Audio-Plus-Vanilla (S1) 1.51 VITA-Audio-Plus-Vanilla 1.94 AISHELL-2 ios AISHELL_2 Qwen2-Audio-base  Qwen2-Audio 3.08 Baichuan-Audio-base  Baichuan-Audio 3.87 Step-Audio-chat  Step-Audio 3.89 Qwen2.5-Omni  Qwen2.5-Omni 2.56 VITA-Audio-Plus-Vanilla 3.29 WenetSpeech WenetSpeech test-meetingtest-net Qwen2-Audio-base  Qwen2-Audio 8.40 — 7.64 Baichuan-Audio-base  Baichuan-Audio 13.28 — 10.13 Freeze-Omni  Freeze-Omni 13.46 — 11.80 VITA-1.5 (S3)  VITA-1.5 10.0 — 8.40 LUCY (S1)  LUCY 10.42 — 8.78 Step-Audio-chat  Step-Audio 10.83 — 9.47 Qwen2.5-Omni  Qwen2.5-Omni 7.71 — 6.04 VITA-Audio-Turbo 23.97 — 18.66 VITA-Audio-Vanilla 17.34 — 13.45 VITA-Audio-Plus-Boost 9.38 — 8.97 VITA-Audio-Plus-Vanilla (S1) 6.68 — 6.59 VITA-Audio-Plus-Vanilla 7.12 — 6.90
Table E5: More results on Spoken Question Answering (SQA) benchmarks. “Sxx” denotes the $x$-th training stage of speech models.
Model #Params Llama Question TriviaQA Web Question Mean S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S Proprietary Models MinMo  MinMo 7B 78.9 64.1 48.3 37.5 55.0 39.9 60.7 47.2 Open-source Models Moshi  Moshi 7B 62.3 21.0 22.8 7.3 26.6 9.2 37.2 12.5 GLM-4-Voice  GLM-4-Voice 9B 64.7 50.7 39.1 26.5 55.0 39.9 45.3 31.0 LUCY (S2)  LUCY 7B 59.6 51.0 23.2 18.2 26.6 18.2 36.5 29.1 MiniCPM-o2.6  yao2024minicpm 7B - 61.0 - 40.0 - 40.2 - 47.0 Llama-Omni  LLaMA-Omni 7B - 45.3 - 22.9 - 10.7 - 26.3 VITA-Audio-Boost 7B 68.7 60.3 30.5 29.3 32.9 30.4 44.0 40.0 VITA-Audio-Vanilla 7B 71.3 66.3 31.9 30.1 33.5 31.4 45.6 42.6 VITA-Audio-Plus-Boost 7B 76.3 64.6 43.6 39.5 44.2 40.0 54.7 48.0 VITA-Audio-Plus-Vanilla 7B 75.6 68.0 45.9 42.7 45.0 41.7 55.5 50.8
Table E6: We have tested VITA-Audio after aligning with ASR and TTS tasks on several text modality benchmarks. After the ASR and TTS alignment, the model’s original text understanding capabilities were indeed affected. However, this might be due to the lack of large-scale, high-quality text data used during training. Interestingly, we observed a slight improvement in performance on GSM8K, which may be due to the fact that our training dataset included a significant amount of math-related data.
Model MMLUhendryckstest2021 GSM8Kcobbe2021gsm8k
Qwen-7B-Instruct 74.22 80.06
VITA-Audio-Plus-Vanilla 66.92 80.14
Table E7: Generation time (ms) of the first audio segment under different inference modes in streaming inference. To enable more real-time speech generation, we progressively increase the number of steps in the flow matching model during streaming inference. The table shows the decoding time when the sampling step of the flow matching model is set to $1$.
VITA-Audio-Boost 39 53 151 243
VITA-Audio-Vanilla 39 236 151 426